# Generative AI and Firm Productivity:

### v1

@@ 5.2 Heterogeneous Effects Across Consumers
4 Marketing Sales 0.6% 7.2% 2.3% 1.7% -6.1% 22.9%***
Push Message Conversion 1.9% 5.5%** 0.5% 5.5%*** -2.3% 15.4%***
@@
    We classify consumers into high and low groups and capture their online shopping experience using three pre-experiment indicators. Consumers in the low group are considered relatively inexperienced, as determined if they meet the following criteria: (1) a registration duration below the median among all consumers on the platform; (2) a number of 30-day login days below the platform-wide median; or (3) a 30-day total purchase amount below the platform-wide median.
@@
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experiment period, and 0 otherwise.
@@
In this experiment, increases in sales and conversion rates were significant and pronounced only for less experienced consumers, while the effects for more experienced consumers were not statistically significant. This pattern likely reflects the challenges faced by inexperienced consumers, who generally possess weaker information-search skills and less familiarity with the online search environment, making it more difficult for them to articulate their needs effectively through query-based searches.
@@
Augmenting product descriptions with GenAI-generated content led to substantial increases in consumer purchases, particularly among inexperienced consumers, as reflected in significant sales gains for this group. A plausible explanation is that inexperienced consumers often lack domain knowledge and access to alternative information sources relative to experienced users, and therefore rely more heavily on detailed descriptions to guide their decisions. These findings suggest that GenAI can help bridge informational gaps and improve the shopping experience for less experienced consumers.
@@
This experiment is not suitable for consumer heterogeneity analysis, due to the fact that the it was conducted off-site at product level, preventing the platform from accessing detailed consumer demographic data from Google.
To sum up, we find that inexperienced consumers—those with shorter registration histories, lower login frequencies, and lower spending levels—benefit more from GenAI than their more experienced counterparts. This result is consistent with prior evidence that GenAI can support vulnerable groups such as low-skilled workers [15, 54, 56], and it extends this insight to the consumer domain by showing that GenAI disproportionately benefits less sophisticated consumers. It also aligns with existing research indicating that enhanced e-commerce technologies—such as improved information provision, search refinement, and personalized recommendation systems—deliver greater benefits to consumers who are older, newer to the platform, or have lower purchasing power [33, 60], while extending these patterns to the context of GenAI.
@@ 5.3 Heterogeneous Effects Across Products
4 Marketing Sales 4.4% -1.5% -0.5% 3.0% 0.6% 2.7%
Push Message Conversion 2.4% 3.5%* 0.6% 5.2%*** 8.9%*** -0.2%
@@
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experiment period, and 0 otherwise.
@@
We find significant gains from GenAI-generated product descriptions in high-concentration categories (Column 1), in high-priced products (Column 5), and across both head and tail products (Columns 3 and 4). These patterns are consistent with several mechanisms. In high-concentration categories such as laptops, product attributes are more standardized and therefore easier to communicate through text-based descriptions. For high-priced items, more comprehensive descriptions may help mitigate information asymmetry, which is particularly salient in costly purchase decisions. The results for head versus tail products may partly reflect the experimental context: the intervention was implemented on platform self-sold products, where Chinese vendors typically provide image-based content that often misaligns with cross-border consumer preferences. Consequently, the head–tail classification may not accurately reflect the quality of text-based description content.
@@
Our results show no significant effects of GenAI deployment on sales across product segments, consistent with the overall analysis. However, we find significant conversion gains in categories with low market concentration (Column 2), in tail products (Column 4), and in high-priced products (Column 5). These patterns align with intuitive mechanisms. In categories such as apparel and fashion, where products are highly differentiated and consumer choices hinge on subtle differences in style, quality, and branding, GenAI-powered marketing—especially when personalized—can more effectively capture attention. Tail products, which typically lack historical data or reputational signals, benefit disproportionately from enhanced marketing outreach. For high-priced items, tailored messaging helps emphasize value, quality assurance, and differentiation, making premium offerings more persuasive.
@@ 6 Discussion and Conclusions
The rapid advances in GenAI have generated widespread expectations among investors and business leaders, driving unprecedented investment in infrastructure and applications. Yet doubts remain regarding the extent to which GenAI can generate substantial productivity improvements at scale. This paper offers some of the first large-scale, real-world evidence on GenAI adoption in online retail, shedding light on how firm-level deployment translates into tangible consumer value and measurable business outcomes.
Our findings yield three main insights. First, GenAI can deliver measurable productivity improvements with comparable inputs, as reflected in increased sales across several business workflows. While the magnitude of these gains varies widely—from negligible effects to double-digit increases—the evidence demonstrates that GenAI can generate substantial economic value when deployed in consumer-facing processes. Second, these improvements arise primarily from enhanced consumer experience and satisfaction through the reduction of frictions in the marketplace, rather than through cost savings on the input side. Across workflows, we observe higher conversion rates: by enriching pre-sale communication, refining search queries, generating richer product descriptions, and personalizing marketing messages, GenAI improves matching efficiency and mitigates information asymmetries. Third, the benefits of GenAI adoption are heterogeneous. Smaller sellers, less experienced consumers, tail products, and high-priced items derive disproportionately larger gains, highlighting GenAI’s role in bridging capability gaps across different segments of the marketplace.
Because most experiments were randomized at the consumer level and overlap across experiments was minimal (less than 1%), the observed effects capture incremental demand (i.e., market expansion) rather than substitution across products. Back-of-the-envelope calculations—annualizing workflow-specific gains and assuming linear additivity—suggest that the four GenAI applications with positive sales effects generate an annual incremental value of approximately \$4.6–\$5 per consumer. These effects represent roughly 5.5–6% of global per-user e-commerce revenue growth in 2023–2024. Taken together, these figures show that even a few GenAI applications already yield substantial gains for a large, mature retailer, with the potential for much larger effects as adoption broadens and increasingly targets revenue-critical workflows. By 2025, the partner platform had deployed GenAI in more than 60 workflows, with usage rising twentyfold as API calls to large language models increased from 2024 to 2025.
@@
Another limitation concerns external validity in general equilibrium. Our experiments were conducted on a single, albeit large, global retail platform. The effects we document partly reflect relative improvements in user experience and satisfaction within this environment. If GenAI adoption becomes widespread and competing platforms deploy similar technologies, these relative advantages may diminish. However, if the mechanism we highlight—enhanced consumer experience through reduced frictions—holds more broadly, there may still be scope for market expansion even in a more competitive environment. In addition, our experiments abstract from strategic responses by competitors —such as changes in pricing or advertising strategy—, which could either amplify or attenuate realized productivity gains.
@@ Appendix A Illustrative User Interfaces and Examples of Each Experiment
In Figure A2, Panel (a) displays the consumer’s original search query in spanish along with the corresponding results, while Panel (b) presents the structured English query translated and refined by GenAI, along with the corresponding search results retrieved. The GenAI-powered query refinement can improve the expression of consumer demand and facilitate the matching efficiency of search engines.
@@ Appendix B Covariance Balance Checks for the Field Experiments
Table B1: Covariate Balance and Randonmization Check
@@
    Mean (Std. Dev.) are shown with all values normalized by the corresponding variable’s mean in the control group. For the definitions of each variable, please refer to the notes in Figure 1.
@@
    The first four experiments are conducted at the consumer level and thus the unit of observation is consumer. The Google Advertising Title experiment is conducted at the product level and thus the unit of observation is product.
@@ Appendix C Main Results: Model, Estimation, and Additional Outcomes
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experiment period, and 0 otherwise.
@@
    “View” stands for the number of product views in the initial search results pages. “Click” stands for the number of product clicks into product detail pages. “Order” is the number of product orders. “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experiment period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase. “Click-through Rate” is the ratio of the number of product clicks to the number of product views. It measures the degree of conversion from views to clicks. “Click-to-Order Conversion Rate” stands for the ratio of the number of product orders to the number of product clicks. It measures the degree of conversion from clicks to purchases.
@@
    “Click” stands for the number of product clicks into product detail pages. “Order” is the number of product orders. “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experiment period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase.
@@
In Table C5, Column 1 reveals that GenAI-generated content accounts for 40% of the marketing push message in the treatment group. With the use of GenAI-generated messages, there are observable rises in consumer clicks and purchases, with the number of clicks increasing by 3.0%, the number of orders increasing by 2.8%, and the purchase amount growing by 1.6% (Columns 2 to 4). Additionally, the likelihood of a consumer making a purchase increased by 3.1% in Column 5. The Cart Value is not statistically significant (Column 6). The results highlight the benefits of using GenAI in unlocking the personalization potential of marketing content, particularly in environments where human-generated content alone may be limited by resource constraints.
@@
    “Is AI Task” is a binary variable which equals one if the message is generated by AI. “Click” stands for the number of clicks on the marketing messages. “Order” is the number of product orders. “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experiment period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase.
@@
where $i$ denotes a product, $y_{i}$ stands for the outcome variables for a product, $Treat_{i}$ is an indicator for whether the product is assigned to the treatment group. We control for product entry-day cohort fixed effect $\alpha_{c{i}}$.
Table C6 presents the main findings. Columns 1 and 2 indicated a 7.6% decrease in ad views and an 10.2% decrease in ad clicks for the treatment group, respectively. Columns 3, 4 and 5 reported non-significant reduction in sales, conversion rate, and cart value, respectively. As we discussed in Section 4.1, the null effect on sales can be attributed to the lack of fine-tuning using e-commerce domain knowledge when setting up the GenAI model.
@@
    “View” represents the number of times the product advertisement is viewed on Google. “Click” refers to the number of times the product advertisement is clicked on Google. “Sales” represents the total expenditure on product orders. “Conversion Rate” measures users’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experiment period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase.
@@
Online sellers often struggle to defend against chargebacks due to reasons such as non-receipt of goods. Chargebacks can lead to significant financial losses and jeopardize the long-term sustainability of sellers’ businesses on the platform. The whole process of contesting chargebacks disputes includes analyzing claims, collecting necessary evidence (e.g., order details, fulfillment records, proof of shipment and logistics tracking through ways like interfacing with diverse APIs), and crafting compelling chargeback defense letters. The rapid advancement of GenAI enabled the platform to develop a chargeback defense agent that offers a one-stop solution to streamlining the intricacies of chargeback disputes.
@@
E-commerce platforms must provide robust consumer services for consumers seeking consultation or negotiation with the platform, such as addressing inquiries about the platform’s promotional details and resolving disputes when consensus with sellers is not reached. For our focal platform, delivering native-language consumer services to a diverse, multilingual consumer base incurs significant costs. Thus, a significant portion of non-English consumer inquires were supported by Filipino consumer service agents in English. A straightforward application of GenAI allows the platform to equip low-cost Filipino agents with robust real-time translation support, aiming to provide more effective communication between consumers and agents.
@@
Appendix D Details on the Heterogeneous Treatment Effects Across Seller

### v2

@@ 5.2 Heterogeneous Effects Across Consumers
4 Marketing Sales -0.9% 6.6% 2.5% 0.8% -5.9% 20.6%***
Push Message Conversion 1.8% 5.0%** 0.7% 4.8%*** -2.2% 14.2%***
@@
    We classify consumers into high and low groups and capture their online shopping experience using three pre-experiment indicators. Consumers in the low group are considered relatively inexperienced, as determined if they meet the following criteria: (1) a registration duration below the platform-wide median; (2) a number of 30-day login days below the platform-wide median; or (3) a 30-day total purchase amount below the platform-wide median.
@@
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise.
@@
In this experiment, increases in sales and conversion rates were significant and pronounced only for less experienced consumers, while the effects for more experienced consumers were not statistically significant. This pattern likely reflects the challenges faced by inexperienced consumers, who generally possess weaker information-search skills and less familiarity with the online search environment. Consequently, they may struggle to articulate their needs effectively through query-based searches and thus derive greater benefits from GenAI-enhanced search.
@@
Augmenting product descriptions with GenAI-generated content led to substantial increases in consumer purchases, particularly among inexperienced consumers, as reflected in the significant sales gains for this group. A plausible explanation is that inexperienced consumers often lack domain knowledge and access to alternative information sources relative to experienced users, and therefore rely more heavily on detailed descriptions to guide their decisions. These findings suggest that GenAI can help bridge informational gaps and improve the shopping experience for less experienced consumers.
@@
This experiment is not suitable for consumer heterogeneity analysis, due to the fact that the it was conducted off-site at the product level, preventing the platform from accessing detailed consumer demographic data from Google.
To sum up, we find that inexperienced consumers—those with shorter registration histories, lower login frequencies, and lower spending levels—benefit more from GenAI than their more experienced counterparts. This result is consistent with prior evidence that GenAI can support vulnerable groups such as low-skilled workers [15, 54, 56], and it extends this insight to the consumer domain by showing that GenAI disproportionately benefits less sophisticated consumers more. It also aligns with existing research indicating that enhanced e-commerce technologies—such as improved information provision, search refinement, and personalized recommendation systems—deliver greater benefits to consumers who are older, newer to the platform, or have lower purchasing power [33, 60], while extending these patterns to the new context of GenAI.
@@ 5.3 Heterogeneous Effects Across Products
4 Marketing Sales -0.8% 5.8% -0.9% 2.4% -5.2% 4.9%
Push Message Conversion 2.9% 3.0%* 0.8% 7.0%*** 4.1%*** 2.7%
@@
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise.
@@
We find significant gains from GenAI-generated product descriptions in high-concentration categories (Column 1), in high-priced products (Column 5), and across both head and tail products (Columns 3 and 4). These patterns are consistent with several mechanisms. In high-concentration categories such as laptops, product attributes are more standardized and therefore easier to communicate through text-based descriptions. For high-priced items, more comprehensive descriptions may help mitigate information asymmetry, which is particularly salient in costly purchase decisions. The results for head versus tail products may partly reflect the experimental context: the intervention was implemented on platform self-sold products, where Chinese vendors typically provide image-based content that often misaligns with cross-border consumer preferences. Consequently, the head–tail classification may not accurately reflect the quality of text-based description content. For instance, a head product may feature high-quality image-based content but still lack a well-developed text description.
@@
Our results show no significant effects of GenAI deployment on sales across product segments, consistent with the main analysis. However, we find significant conversion gains in categories with low market concentration (Column 2), in tail products (Column 4), and in high-priced products (Column 5). These patterns align with intuitive mechanisms. In categories such as apparel and fashion, where products are highly differentiated and consumer choices hinge on subtle differences in style, quality, and branding, GenAI-powered marketing—especially when personalized—can more effectively capture attention. Tail products, which typically lack historical data or reputational signals, also benefit more from enhanced marketing outreach. For high-priced items, tailored messaging helps emphasize value, quality assurance, and differentiation, making premium offerings more persuasive.
@@ 6 Discussion and Conclusions
The rapid advances in GenAI have generated widespread expectations among investors and business leaders, driving unprecedented investment in infrastructure and applications. Yet doubts remain regarding the extent to which GenAI can generate substantial productivity improvements at scale. This paper offers some of the first large-scale, real-world experimental evidence on GenAI adoption in online retail, shedding light on how firm-level deployment translates into tangible consumer value and measurable business outcomes.
Our findings yield three main insights. First, GenAI can deliver measurable productivity improvements with comparable inputs, as reflected in increased sales across several business workflows. While the magnitude of these gains varies widely—from negligible effects to double-digit increases—the evidence demonstrates that GenAI can generate substantial economic value when deployed in consumer-facing processes. Second, these improvements arise primarily from enhanced consumer experience and satisfaction through the GenAI-driven reduction of frictions in the marketplace, rather than from cost savings on the input side. Across workflows, we observe higher conversion rates: by enriching pre-sale communication, refining search queries, generating richer product descriptions, and personalizing marketing messages, GenAI improves matching efficiency and mitigates information asymmetries. Third, the benefits of GenAI adoption are heterogeneous. Smaller sellers, less experienced consumers, tail products, and high-priced items derive disproportionately larger gains, highlighting GenAI’s role in bridging capability gaps across different segments of the marketplace.
Because most experiments were randomized at the consumer level and overlap across experiments was minimal (less than 1%), the observed effects capture incremental demand (i.e., market expansion) rather than substitution across products. Back-of-the-envelope calculations—annualizing workflow-specific gains and assuming linear additivity—suggest that the four GenAI applications with positive sales effects generate an annual incremental value of approximately \$4.6–\$5 per consumer. These effects represent roughly 5.5–6% of global per-user e-commerce revenue growth in 2023–2024. Taken together, these figures show that even a few GenAI applications already yield substantial gains for a large, mature retailer, with the potential for much larger effects as adoption broadens and increasingly targets revenue-critical workflows. By 2025, the partner platform had deployed GenAI in more than 60 workflows, with usage rising twentyfold as API calls to its proprietary GenAI tools increased between 2024 and 2025.
@@
Another limitation concerns external validity in general equilibrium. Our experiments were conducted on a single, albeit large, global retail platform. The effects we document partly reflect relative improvements in user experience and satisfaction within this environment. If GenAI adoption becomes widespread and competing platforms deploy similar technologies, these relative advantages may diminish. However, if the mechanism we highlight—enhanced consumer experience through reduced frictions—holds more broadly, there may still be scope for market expansion even in a more competitive environment. In addition, our experiments abstract from strategic responses by competitors —such as changes in pricing or advertising strategy— which could either amplify or attenuate realized productivity gains.
@@ Appendix A Illustrative User Interfaces and Examples of Each Experiment
In Figure A2, Panel (a) displays the consumer’s original search query in Spanish, along with the corresponding results, while Panel (b) presents the structured English query translated and refined by GenAI, as well as the corresponding search results retrieved. The GenAI-powered query refinement can improve the expression of consumer demand and facilitate the matching efficiency of search engines.
@@ Appendix B Covariance Balance Checks for the Field Experiments
Table B1: Covariate Balance and Randomization Checks
@@
    Mean (Std. Dev.) are shown with all values normalized by the corresponding variable’s control group mean. For the definitions of each variable, please refer to the notes in Figure 1.
@@
    The first four experiments are conducted at the consumer level and thus the unit of observation is the consumer. The Google Advertising Title experiment is conducted at the product level and thus the unit of observation is the product.
@@ Appendix C Main Results: Model, Estimation, and Additional Outcomes
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise.
@@
    “View” stands for the number of product views in the search results pages. “Click” stands for the number of product clicks into product detail pages. “Order” is the number of product orders. “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase. “Click-through Rate” is the ratio of the number of product clicks to the number of product views. It measures the degree of conversion from views to clicks. “Click-to-Order Conversion Rate” stands for the ratio of the number of product orders to the number of product clicks. It measures the degree of conversion from clicks to purchases.
@@
    “Click” stands for the number of product clicks into product detail pages. “Order” is the number of product orders. “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase.
@@
In Table C5, Column 1 reveals that GenAI-generated content accounts for 40% of the marketing push message in the treatment group. With the use of GenAI-generated messages, there are observable rises in consumer clicks and purchases, with the number of clicks increasing by 3.1%, the number of orders increasing by 2.8%, and the purchase amount growing by 1.6% (Columns 2 to 4). Additionally, the likelihood of a consumer making a purchase increased by 3% in Column 5. The Cart Value is not statistically significant (Column 6). The results highlight the benefits of using GenAI in unlocking the personalization potential of marketing content, particularly in environments where human-generated content alone may be limited by resource constraints.
@@
    “Is AI Task” is a binary variable which equals one if the message is generated by AI. “Click” stands for the number of clicks on the marketing messages. “Order” is the number of product orders. “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase.
@@
where $i$ denotes a product, $y_{i}$ stands for the outcome variables for a product, $Treat_{i}$ is an indicator for whether the product is assigned to the treatment group. We control for product entry-day cohort fixed effect $\alpha_{c(i)}$.
Table C6 presents the main findings. Columns 1 and 2 indicated a 7.6% decrease in ad views and a 10.2% decrease in ad clicks for the treatment group, respectively. Columns 3, 4, and 5 reported a non-significant reduction in sales, conversion rate, and cart value, respectively. As we discussed in Section 4.1, the null effect on sales can be attributed to the lack of fine-tuning using e-commerce domain knowledge when setting up the GenAI model.
@@
    “View” represents the number of times the product advertisement is viewed on Google. “Click” refers to the number of times the product advertisement is clicked on Google. “Sales” represents the total expenditure on product orders. “Conversion Rate” measures users’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise. “Cart Value” refers to the expenditure per product, conditional on the product being purchased.
@@
Online sellers often struggle to defend against chargebacks due to reasons such as non-receipt of goods. Chargebacks can lead to significant financial losses and jeopardize the long-term sustainability of sellers’ businesses on the platform. The whole process of contesting chargeback disputes includes analyzing claims, collecting necessary evidence (e.g., order details, fulfillment records, proof of shipment and logistics tracking through ways like interfacing with diverse APIs), and crafting compelling chargeback defense letters. The rapid advancement of GenAI enabled the platform to develop a chargeback defense agent that offers a one-stop solution to streamlining the intricacies of chargeback disputes.
@@
E-commerce platforms must provide robust consumer services for consumers seeking consultation or negotiation with the platform, such as addressing inquiries about the platform’s promotional details and resolving disputes when consensus with sellers is not reached. For our focal platform, delivering native-language consumer services to a diverse, multilingual consumer base incurs significant costs. Thus, a significant portion of non-English consumer inquiries were supported by Filipino consumer service agents in English. A straightforward application of GenAI allows the platform to equip low-cost Filipino agents with robust real-time translation support, aiming to provide more effective communication between consumers and agents.
@@
Appendix D Details on the Heterogeneous Treatment Effects Across Sellers